BOUNDS ON VOLUME INCREASE

UNDER DEHN DRILLING OPERATIONS

Martin Bridgeman

Abstract. In this paper we investigate how the volume of hyperbolic manifolds increases under the process of removing a curve, that is, Dehn drilling. If the curve we remove is a geodesic we are able to show that for a certain family of manifolds the volume increase is bounded above by π · l where l is the length of the geodesic drilled. Also we construct examples to show that there is no lower bound to the volume increase in terms of a linear function of a positive power of length and in particular volume increase is not bounded linearly in length.

Typeset by AMS-TEX